OpenStudy (anonymous):

L'Hopitals' Rule - can anyone check my answer? lim x-> inf of (x^2) / (ln x) inf / inf form, so lim x->inf of 2x^2 and lim x-> inf of 2(inf)^2 = inf <--final answer. Is this right?

OpenStudy (anonymous):

it is in that form, but you should do it in your head

OpenStudy (anonymous):

log grows very very very slowly, much slower than any polynomial

OpenStudy (anonymous):

I should do it in my head?

OpenStudy (anonymous):

i don't understand.. so where did i get it wrong?

OpenStudy (anonymous):

but yes, if you use l'hopital you get the right answer i just mean you should know in advance what the answer is

OpenStudy (anonymous):

you didn't do it wrong, you did it right

OpenStudy (anonymous):

oh... well but i don't know... haha

OpenStudy (anonymous):

log grows slower than \(x\) or \(x^2\) so \[\lim_{x\to \infty}\frac{x^2}{\log(x)}=\infty\] for sure

OpenStudy (anonymous):

think if it was base ten (which it isn't but might as well be) if you had \(x=1,000,000\) then it would be \[\frac{1,000,000,000,000}{6}\] a pretty big number!

OpenStudy (anonymous):

i dont think i get infinite limits in the first place... wow...

OpenStudy (anonymous):

so i''m ust following the step by step directions for finding these limits with l'hopital rule...

OpenStudy (anonymous):

Thank you again @satellite73

OpenStudy (anonymous):

you will develop intuition as you go along

OpenStudy (anonymous):

thank you @thomasz ! but i did the same work you did to find the answer :)

OpenStudy (anonymous):

You too! :D @thomasz